OpenStudy (anonymous):

Find the length of the curve: x=6cos(t), y=6sin(t), 0 <= t <= pi

OpenStudy (anonymous):

Not really sure how to go about this, I gather it has something to do with polar coordinates but I'm not really sure how to convert that into a polar equation.

OpenStudy (anonymous):

you just need to apply the formula for the arc length of the curve with parametric equations\[L = \int\limits_{a}^{b}\sqrt{\left( \frac{dx}{dt} \right)^2 + \left( \frac{dy}{dt} \right)^2}dt\]

OpenStudy (anonymous):

oh thanks

OpenStudy (amistre64):

orrr, just find the circumference of half a circle with a radius of 6